# Test-Time Adaptation via Many-Shot Prompting: Benefits, Limits, and Pitfalls

### 12 Mar 2026 v2

@@ 4.1 More Context Helps—Until It Doesn’t
Unless otherwise stated, we refer to $n$ as the per-class shot count and $N$ as the total number of demonstrations. As shown in Fig. 2(a), as the update magnitude increases for LLaMA-3.1-8B-Instruct, accuracy improves steadily, after which it reaches a plateau at approximately 50–70 shots per class. Beyond this point, additional demonstrations provide diminishing returns. This behavior indicates a clear saturation regime: moderate input-space updates introduce useful task-specific signal, while aggressive updates yield diminishing further gains Agarwal et al. (2024). We hypothesize that the saturation observed in many-shot prompting arises from the limited representational capacity of attention-based conditioning. Demonstrations influence predictions through their key–value representations in the context, but the resulting update is a weighted aggregation. Once sufficient demonstrations are present to identify the task pattern, additional examples contribute largely redundant representations, causing marginal gains to diminish and, in very long prompts, attention competition can even lead to performance degradation.
@@ 4.2 Update policy matters—relevance helps early, diversity helps at scale
Next, we study how update policy affects test-time adaptation under a fixed context budget in Fig. 2(b) for LLaMA-3.1-8B-Instruct. Across all comparisons, we fix the total number of demonstrations $N$ and vary only the update policy used to select these. We ensure prompt template consistency and prevent including query or same-text duplicates in selected examples.
Label-wise vs. Cross-label: We construct prompts using a per-class formulation with $n$ demonstrations per class, resulting in $N=n\times C$ total demonstrations for Banking77. In label-wise selection, we enforce balanced coverage by selecting exactly $n$ demonstrations per class. In cross-label selection, we select $N$ demonstrations from the entire dataset, so label frequencies can be uneven.
@@
Comparing these policies, cross-label selection consistently outperforms label-wise selection, suggesting that enforcing per-label balance can reduce useful diversity by over representing redundant examples. Cross-label benefits from exposure to greater contextual diversity thereby improving generalization, label-balance forces many low information examples which is not very helpful. Cross-label similarity-based selection is strong at small update magnitudes (high relevance), but degrades as $N$ grows, whereas cross-label random selection scales more robustly with larger $N$ (higher diversity). This reflects a bias-variance tradeoff in test-time updates: relevance-focused (similarity-based) policies help early but can over concentrate the context as the update becomes aggressive; diversity-focused (random) policies scale better. Our best setting is ($n=1\Rightarrow N=77$) for cross label similarity. This maximizes relevance per demonstration per class label, yielding strong task adaptation before redundancy degrades performance at larger scale. Since cross-label sampling does not enforce uniform class coverage, improvements in overall accuracy may partially reflect gains on frequent classes; future work could analyze per-class accuracy or macro-F1 to better understand the impact on tail classes. Additional results included in Appendix A.3.
@@ A.2 Additional Experimental Details for Update Magnitude Experiments
For the results shown in Fig. 2(a) on Banking77, we vary the per-class shot count $n$. The total number of demonstrations is given by $N=n\times C$, where $C=77$ is the number of classes. Table 1 illustrates how $N$ scales with $n$. We stop at $N=5390$ demonstrations, corresponding to the maximum supported context length of 128k tokens for LLaMA models.
@@ A.3 Dynamic ICL selection strategies
  • Cross-label: select $N$ demonstrations globally; class counts are unconstrained.
@@
  • Similarity: top-$N$ retrieval by embedding similarity to the query.

### 17 Mar 2026 v3

@@ 4.1 More Context Helps—Until It Doesn’t
Unless otherwise stated, we refer to $n$ as the per-class shot count and NN as the total number of demonstrations. As shown in Fig. 2(a), as the update magnitude increases for LLaMA-3.1-8B-Instruct, accuracy improves steadily, after which it reaches a plateau at approximately 50–70 shots per class. Beyond this point, additional demonstrations provide diminishing returns. This behavior indicates a clear saturation regime: moderate input-space updates introduce useful task-specific signal, while aggressive updates yield diminishing further gains Agarwal et al. (2024). We hypothesize that the saturation observed in many-shot prompting arises from the limited representational capacity of attention-based conditioning. Demonstrations influence predictions through their key–value representations in the context, but the resulting update is a weighted aggregation. Once sufficient demonstrations are present to identify the task pattern, additional examples contribute largely redundant representations, causing marginal gains to diminish and, in very long prompts, attention competition can even lead to performance degradation.
@@ 4.2 Update policy matters—relevance helps early, diversity helps at scale
Next, we study how update policy affects test-time adaptation under a fixed context budget in Fig. 2(b) for LLaMA-3.1-8B-Instruct. Across all comparisons, we fix the total number of demonstrations NN and vary only the update policy used to select these. We ensure prompt template consistency and prevent including query or same-text duplicates in selected examples.
Label-wise vs. Cross-label: We construct prompts using a per-class formulation with $n$ demonstrations per class, resulting in $N=n\times C$ total demonstrations for Banking77. In label-wise selection, we enforce balanced coverage by selecting exactly $n$ demonstrations per class. In cross-label selection, we select NN demonstrations from the entire dataset, so label frequencies can be uneven.
@@
Comparing these policies, cross-label selection consistently outperforms label-wise selection, suggesting that enforcing per-label balance can reduce useful diversity by over representing redundant examples. Cross-label benefits from exposure to greater contextual diversity thereby improving generalization, label-balance forces many low information examples which is not very helpful. Cross-label similarity-based selection is strong at small update magnitudes (high relevance), but degrades as NN grows, whereas cross-label random selection scales more robustly with larger NN (higher diversity). This reflects a bias-variance tradeoff in test-time updates: relevance-focused (similarity-based) policies help early but can over concentrate the context as the update becomes aggressive; diversity-focused (random) policies scale better. Our best setting is ($n=1\Rightarrow N=77$) for cross label similarity. This maximizes relevance per demonstration per class label, yielding strong task adaptation before redundancy degrades performance at larger scale. Since cross-label sampling does not enforce uniform class coverage, improvements in overall accuracy may partially reflect gains on frequent classes; future work could analyze per-class accuracy or macro-F1 to better understand the impact on tail classes. Additional results included in Appendix A.3.
@@ A.2 Additional Experimental Details for Update Magnitude Experiments
For the results shown in Fig. 2(a) on Banking77, we vary the per-class shot count $n$. The total number of demonstrations is given by $N=n\times C$, where $C=77$ is the number of classes. Table 1 illustrates how NN scales with $n$. We stop at $N=5390$ demonstrations, corresponding to the maximum supported context length of 128k tokens for LLaMA models.
@@ A.3 Dynamic ICL selection strategies
  • Cross-label: select NN demonstrations globally; class counts are unconstrained.
@@
  • Similarity: top-NN retrieval by embedding similarity to the query.